Meeting Notes for 18 May 2011

Present: Phong, TomL, TomJ, Liz
Absent: Marge, Bob, TomM


  • Compute nodes upgraded and operational
    • node7 froze sometime Monday night, investigating (and running stress tests)
  • New data server arrived; not yet in production (undergoing testing)
  • Server information up at


  • Blog searching (across categories, tags, etc)


  • Blessing marching forward
  • Someone hammered www18 yesterday but it stayed up without flinching


  • List of things to do was sent to the mailing list
  • How to deal with muon (and electron) quality cuts?

-- Main.PhongNguyen - 18 May 2011
Topic revision: r15 - 2011-05-18, PhongNguyen
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback